# Hyper-STTN: Social Group-aware Spatial-Temporal Transformer Network for Human Trajectory Prediction with Hypergraph Reasoning

### v2

@@ III-D Hypergraph Neural Network
The crowd groups of Hyper-STTN are constructed with respect to neighborhoods’ feature dimensional Mahalonobis distance [21], considering both interactive correlation distribution and physical spatial distance. In particular, the pedestrians’ motion similarity is captured by the covariance matrix of Mahalonobis distance to cover the situations that a long-distance person presents a high potentiality to join the group. We construct crowd hypergraphs by vertex classifications of social group which are formulated as a spectral hypergraph k-way partitioning problem [22, 23]. The normalized one cut spectral hypergraph partitioning task is defined as follows:
@@
Assuming the $k$-way hypergraph partition by a set of vertexes subsets $\{\mathcal{V}_{1},\cdots,\mathcal{V}_{p}\}$ from $\mathbf{F}=[f_{1}\cdots f_{p}]$. Subsequently, the spectral hypergraph $k$-way partitioning problem as a combinatorial optimization problem can be relaxed with respect to minimizing $\beta(\mathcal{V}_{1},\cdots,\mathcal{V}_{p})$ as follows:
@@
To address above k-way hypergraph partition task, the K-nearest neighbor (KNN) [24] method is leveraged for multiscale hypergraphs generation. Hyper-STTN iterates each vertex who presents a person with KNN algorithm to search several its interactive neighborhoods based on Mahalonobis feature distance. In details, the low-level trajectory embeddings of each agent $\{\mathbf{q}_{(\rm{X}_{\emph{1}})},\cdots,\mathbf{q}_{(\rm{X}_{\emph{N}})};\mathbf{q}({x}_{i})\in{\mathbb{R}}^{d}\}$ is encoded by a fully connected layer (FC) from input $\{\mathbf{X}_{\emph{1}},\cdots,\mathbf{X}_{\emph{N}}\}$. Subsequently, the similarity matrix $\mathcal{S}$ is filled by feature dimensional agent pairs’ Mahalanobis distance [21], coupling the interaction attributes and motion distributions to dilute Euclidean distance analogous interferences and errors from feature correlations and distributions. The Mahalanobis distance of $(i,j)$-th vertex pair $Dis(i,j)$ in the feature dimension is defined as follows:
@@
Inspired by [25, 23], KNN searches each vertex and its $K$ nearest neighbors to construct multiscale hypergraphs with respect to the similarity matrix $\mathcal{S}$, as shown in Fig. 3. Each vertex of hypergraph represents single agent, and hyperedges could link multiple agents to form group-wise interactions.
After creating crowd multiscale hypergraphs, a hypergraph convolution neural network [23] is introduced to estimate the group-wise interactions. The random walk probability [22] of hypergraph aggregates the weighted dependencies of all sub-vertexes as hyperedges features, and then gathers the weighted dependencies of all hyperedges features together as vertex features as shown in Fig. 3. Assume a vertex $v_{i}$ is stochastically selected with one of its hyperedges $e_{o}$, and the probability of hypergraph random walk $\mathcal{O}_{(v_{i},v_{j})}$ from $v_{i}$ to $v_{j}$ on hyperedge $e_{o}$ can be defined as follows:
@@
Figure 4: Hybrid Spatial-Temporal Transformer Framework: Pedestrians’ motion intents and dependencies are abstracted as spatial and temporal attention maps by multi-head attention mechanism of spatial-temporal transformer. Additionally, a multi-head cross attention mechanism is employed to align heterogeneous group-wise and pair-wise features.
Upon the definition of random walk probability, the hypergraph interaction feature, as group-wise interaction information, is calculated from hypergraph convolution operation based on hypergraph random walk theory, which can aggregate overall dependencies of vertexes and hyperedges on hypergraphs. According to [26, 23], the spectral hypergraph convolution operation that a filter $\mathbf{g}$ spectral convoluted input data $\mathbf{x}\in\{\mathbf{x}_{1},\cdots,\mathbf{x}_{N}\}$ can be defined as follows:
@@
where $\Delta=\mathbf{I}-\mathcal{O}$ is the regularized hypergraph Laplacian matrix. And $\mathbf{T}_{u}(\widetilde{\Delta})$ is the Chebyshev polynomial of order $u$ with scaled Laplacian $\widetilde{\Delta}=2{\Delta}/{\lambda_{\max}}-\mathbf{I}$, where ${\lambda_{\max}}$ is the largest eigenvalue of $\Delta$. The parameter $\theta$ is a weighted parameter, and ${U}$ is the kernel size of the graph convolution.
Decomposing the hypergraph Laplacian matrix and approximating it by the first-order Chebyshev polynomial as a hypergraph convolution operation with single scale group-wise interaction dependence as output $\mathbf{Y}\in\mathbb{R}^{N\times d}$.
@@
Finally, we aggregate all the multiscale hypergraphs together to present crowded miscellaneous group-wise interactions as follows:
@@
where $f_{MLP}$ is a MLP neural network, and $\mathbf{\hat{Y}_{\rm{H}}}\in\mathbb{R}^{N\times H\times d}$ is the group-wise interaction feature.
@@ III-E Multi-Modal Fusion Network and CVAE Decoder
Subsequently, the group-wise HHI features $\mathbf{\hat{Y}}_{\rm{HS}},\mathbf{\hat{Y}}_{\rm{HT}}$ and pair-wise HHI dependencies $\mathbf{\hat{Y}}_{\rm{ST}}$ are aligned by a multi-head self-attention network as final crowd dynamics representation $\mathbf{\hat{Y}}_{\rm{M}}$.
@@
To estimate the stochasticity of human movements, a conditional variational auto encoder (CVAE)-based decoder [18] is employed to approximate maximum likelihood in potential distributions of motion uncertainty. The environmental dynamics feature $\mathbf{\hat{Y}}_{\rm M}$ and observed data $\mathbf{X}$ are encoded to present Gaussian distribution ${\mathcal{N}(\mu,\sigma^{2})}$. And the latent variable $z$ is sampled by above Gaussian distribution $z\sim{\mathcal{N}(0,\sigma_{\rm T_{i}}^{2}\mathbf{I})}$ in the testing process. Lastly, the maximum likelihood forecasting result is calculated by the decoder based on the concatenation of latent variable and observation embedding in the CVAE block as follows:
@@
where $\mu,\sigma$ are the mean and variation of the approximate distribution, and $Encoder(\cdot)$ is Hyper-STTN backbone framework, and $Decoder(\cdot)$ is a spectral temporal graph network from [27].
@@ IV-A1 Dataset
We have conducted experiments on several commonly used public pedestrian trajectory datasets collected from real-world scenarios. The training and validation datasets are consisted of the ETH-UCY [28], NBA, and SDD [29] datasets. These datasets record the temporal location information of pedestrians in world coordinates, each identified by an individual index. Additionally, the experiments of ETH-UCY are conducted on the leave-one-out cross-validation evaluation strategy, where the last one subsection is reserved for testing in each experiment.
@@ IV-A2 Evaluation Metrics
In our experiments, we employed the evaluation metrics Average Displacement Error (ADE) and Final Displacement Error (FDE) [4] to assess prediction accuracy. The ADE/FDE and ADE20/FDE20 metrics measure prediction performance by calculating the Euclidean distance between the prediction and ground truth, either on average or at a single endpoint.
@@ IV-B1 Baselines
We compared our model with several existing SOTA algorithms. The algorithms included in the comparison are Social-Attention [13], Social-GAN [6], Social-STGCNN [15], Trajectron++ [9], STAR [7], PECNet [30], AgentFormer [8], Social-Implicit [31], GroupNet [5], and EqMotion [32]. We trained and tested all baseline networks using the same training datasets and validation datasets obtained from public sources, and we followed the parameters and configurations outlined in their original papers.
@@ IV-B2 Ablation Models
We designed three ablation models to investigate the influence of different subsection blocks on performance. First, STTN is designed to infer underlying pedestrian interactions based solely on the spatial-temporal transformer, without the hypergraph network block from Hyper-STTN. Meanwhile, HGNN is configured to model crowded motion dependencies using a set of multiscale hypergraph constructions and hypergraph convolution operations. The final ablation model (Hyper-STTN+MLP) removes the multimodal transformer block and employs a MLP (multilayer perceptron) for group-wise and pair-wise feature fusion. Additionally, the CVAE structure serves as a decoder for each ablation model. And training configurations, such as the training dataset and initialized parameters, are shared among the ablations models and Hyper-STTN.
@@
Stochastic ADE20 / FDE20 Social- Attention [13] ICRA18 Social- GAN [6] CVPR18 Trajectron ++ [9] ECCV20 STAR [7] ECCV20 Agent- Former [8] ICCV21 Social- Implicit [31] ECCV22 GroupNet [5] CVPR22 EqMotion [32] CVPR23 STTN (ablation) HGNN (ablation) Hyper-STTN +MLP (ablation) Hyper-STTN (ours)
ETH 1.39/2.39 0.87/1.62 0.61/1.02 0.36/0.65 0.45/0.75 0.61/1.08 0.46/0.73 0.40/0.61 0.35/0.60 0.52/0.75 0.42/0.73 0.35/0.59
HOTEL 2.51/2.91 0.67/1.37 0.19/0.28 0.17/0.36 0.14/0.22 0.33/0.63 0.15/0.25 0.12/0.18 0.26/0.50 0.21/0.25 0.21/0.32 0.14/0.19
UNIV 1.25/2.54 0.76/1.52 0.30/0.54 0.31/0.62 0.25/0.45 0.52/1.11 0.26/0.49 0.23/0.43 0.25/0.55 0.25/0.43 0.30/0.53 0.23/0.42
ZARA1 1.01/2.17 0.35/0.68 0.24/0.42 0.26/0.55 0.18/0.30 0.32/0.66 0.21/0.39 0.18/0.32 0.21/0.50 0.31/0.39 0.28/0.53 0.18/0.29
ZARA2 0.88/1.75 0.42/0.84 0.18/0.32 0.22/0.46 0.14/0.24 0.43/0.85 0.17/0.33 0.13/0.23 0.19/0.40 0.25/0.33 0.23/0.28 0.15/0.23
AVG 1.41/2.35 0.61/1.21 0.30/0.51 0.26/0.53 0.23/0.39 0.33/0.67 0.25/0.44 0.21/0.35 0.25/0.51 0.31/0.43 0.29/0.48 0.21/0.34
TABLE II: The crowded trajectory forecasting best-of-20 stochastic sampled results of minADE20 / minFDE20 on NBA &\&& SDD dataset.
Stochastic ADE20 / FDE20 Social-GAN [6] CVPR18 Social-STGCNN [15] CVPR20 Trajectron++ [9] ECCV20 PECNet [30] ECCV20 GroupNet [5] CVPR22 STTN (ablation) HGNN (ablation) Hyper-STTN+MLP (ablation) Hyper-STTN (ours)
NBA-1s 0.41/0.62 0.34/0.48 0.30/0.38 0.35/0.58 0.34/0.48 0.39/0.58 0.37/0.59 0.35/0.48 0.30/0.37
NBA-2s 0.81/1.32 0.71/0.94 0.59/0.82 0.68/1.23 0.62/0.95 0.72/1.05 0.64/0.95 0.65/0.89 0.58/0.81
NBA-3s 1.19/1.94 1.09/1.77 0.85/1.24 1.01/1.76 0.87/1.31 1.07/1.51 0.98/1.44 0.95/1.51 0.84/1.21
NBA-4s 1.59/2.41 1.53/2.26 1.15/1.57 1.31/1.79 1.13/1.69 1.24/1.99 1.18/1.89 1.21/1.75 1.01/1.52
SDD 27.23/41.44 20.8/33.2 19.30/32.70 9.96/15.88 9.31/16.11 11.17/19.28 11.13/18.43 10.11/18.41 9.05/15.14
@@ IV-B3 Training Details
To train the model, we use the Adam optimizer with a learning rate $1\times 10^{-4}$ and a decay rate (50% per 100 epochs). Additionally, Gaussian noise is added to enforce the robustness of the network. For more parameters and details, please refer to our project website.
@@
where $\kappa$ is the weight of each loss term, and the Kullback-Leibler (KL) divergence term ${f_{\mathrm{KL}}(\cdot||\cdot)}$ is used to update the encoder and decoder in the CVAE block, and ${\mathbb{{\hat{X}}}}$ denotes the groundturth of dataset. The ($\angle$($\cdot$)∠ ( ⋅ ) function computes the angle between two vectors representing points.
@@
IV-C Qualitative Results
We conducted the trajectory prediction task with respect to stochastic conditions, which generate 20 potential paths to evaluate with the best one of total samples. As shown in Tables I and II, Hyper-STTN is compared with several baselines and three ablation models on the ETH-UCY dataset, NBA dataset and SDD dataset. Hyper-STTN demonstrates the best performance among current SOTA frameworks, both in ETH-UCY, NBA, and SDD configurations. For instance, in the ETH-UCY dataset, Hyper-STTN not only enhances the $12.5\%$ ADE20 and $3.2\%$ FDE20 from EqMotion’s [32] 0.40/0.61 to 0.35/0.59 on ETH dataset, but also improves the average $14.7\%/6.5\%/3.4\%/10.6\%$ ADE20 and $22.9\%/14.7\%/7.6\%/10.1\%$ FDE20 performance from Groupnet’s [5] on NBA dataset. Highlighting the effectiveness of Hyper-STTN in crowd latent interaction inference and trajectory forecasting.
The insights of deep learning-inspired trajectory prediction framework have motivated the research paradigm that implicitly capture latent social interaction among spatial and temporal dimensions to estimate pedestrian potential movements. Whereas, the deployments of LSTM [4], Attention [13], to transformer [7] have increased the average performance of pedestrians’ trajectory predictors with respect to the more accurate high-level feature inference on advanced framework. Thus, the transformer-based backbones are leveraged by Hyper-STTN to address fundamental social interactions and temporal dependencies among crowded movements. More specifically, the ablation model STTN, which is solely driven by a transformer like STAR. From our experiments, we observed that the aforementioned enrichment is a result of the introduction of masked attention and cross-modal attention mechanisms. These mechanisms not only mitigate the influence of agents flicker when compared to vanilla self-attention in the fixed pre-defined transformer data scale formulation but also enhance the adaptability for spatial-temporal fusion.
However, despite transformer demonstrates the amazing effectiveness on pair-wise interaction abstraction, the overlook of group-wise interaction limits the further performance enhancements, particularly in dense scenarios individual motion is more sensitive for the movements of crowded groups. Such as GroupNet [5] ($0.26/0.49$) is better than STAR [7] ($0.31/0.62$) on UNIV dataset that owns more dense crowded scenarios. Therefore, herein we leverage hypergraph construction on group-wise interaction inference and transformer on pair-wise feature understanding to address crowded movement strategies. Apart from that, the ablation model HGNN exhibits a similar performance with GroupNet on several datasets.

### v3

@@ III-D Hypergraph Neural Network
The crowd groups of Hyper-STTN are constructed with respect to neighborhoods’ feature dimensional Mahalonobis distance [25], considering both interactive correlation distribution and physical spatial distance. In particular, the pedestrians’ motion similarity is captured by the covariance matrix of Mahalonobis distance to cover the situations that a long-distance person presents a high potentiality to join the group. We construct crowd hypergraphs by vertex classifications of social group which are formulated as a spectral hypergraph $k$-way partitioning problem [26, 27]. We define a group at time $t$ as a set of agents whose latent motion embeddings are mutually similar under a Mahalanobis metric computed over short-horizon kinematics (position, velocity, heading) and context crowd distribution features; proximity alone is not sufficient. Hyperedges connect an agent to its top-$k$ neighbors under this metric; varying $k_{i}$ yields multiscale groups with respect to $\tau_{i}$. The normalized one cut spectral hypergraph partitioning task is defined as follows:
@@
Assuming the $k$-way hypergraph partition by a set of vertices subsets $\{\mathcal{V}_{1},\cdots,\mathcal{V}_{p}\}$ from $\mathbf{F}=[f_{1}\cdots f_{p}]$. Subsequently, the spectral hypergraph $k$-way partitioning problem as a combinatorial optimization problem can be relaxed with respect to minimizing $\beta(\mathcal{V}_{1},\cdots,\mathcal{V}_{p})$ as follows:
@@
To address above k-way hypergraph partition task, the K-nearest neighbor (KNN) [28] method is leveraged for multiscale hypergraphs generation. Hyper-STTN iteratively applies KNN to each vertex to identify its interactive neighborhoods based on the Mahalanobis feature distance. Particularly, the hypergraph at scale $\tau_{1}=2$ degenerates to a standard graph, representing only pairwise interaction. In details, the low-level trajectory embeddings of each agent $\{\mathbf{q}_{(\rm{X}_{\emph{1}})},\cdots,\mathbf{q}_{(\rm{X}_{\emph{N}})}\}$, where $\mathbf{q}({x}_{i})\in{\mathbb{R}}^{d}$ are obtained via a fully connected layer from the input sequences $\{\mathbf{X}_{\emph{1}},\cdots,\mathbf{X}_{\emph{N}}\}$. The similarity matrix $\mathcal{S}$ is then filled by feature dimensional agent pairs’ Mahalanobis distance [25], which incorporate interaction attributes and motion distributions to mitigate Euclidean distance based distortions arising from feature correlations and heterogeneous distributions. The Mahalanobis distance of $(i,j)$-th vertex pair $Dis(i,j)$ is defined as follows:
@@
Inspired by [29, 27], we initialize multiscale hypergraphs by performing KNN searches for each vertex and its $k$ nearest neighbors with respect to the similarity matrix $\mathcal{S}$, as shown in Fig. 3. In this representation, each vertex corresponds to a single agent, while hyperedges link multiple agents for the representation of groupwise interaction. The $k$-nearest neighbor selection is kept fixed per training epoch from the current embeddings (stop-gradient on neighbor indices). We refresh neighbor sets every $E$ epochs to account for updated embedding structures during training.
After constructing multiscale crowd hypergraphs, a hypergraph convolution neural network [27] is employed to estimate the groupwise interactions. The random walk probability [26] of hypergraph aggregates the weighted dependencies of all sub-vertices into hyperedges features, which are then combined into vertex features, as shown in Fig. 3. Let a vertex $v_{i}$ be stochastically selected with one of its hyperedges $e_{o}$, and the probability of hypergraph random walk $\mathcal{O}_{(v_{i},v_{j})}$ from $v_{i}$ to $v_{j}$ on hyperedge $e_{o}$ can be defined as follows:
@@
Figure 4: Hybrid Spatial-Temporal Transformer Framework: Pedestrians’ motion intents and dependencies are abstracted as spatial and temporal attention maps by multi-head attention mechanism of spatial-temporal transformer. Additionally, a multi-head cross attention mechanism is employed to align heterogeneous groupwise and pairwise features.
Based on the definition of random walk probability, the hypergraph interaction features, representing groupwise interaction information are calculated by a hypergraph spectral convolution operation grounded, which aggregates the global dependencies between vertices and hyperedges. Following [30, 27], the spectral hypergraph convolution of an input signal $\mathbf{x}\in\{\mathbf{x}_{1},\cdots,\mathbf{x}_{N}\}$ with a filter $\mathbf{g}$ is defined as follows:
@@
[…]
Decomposing the hypergraph Laplacian matrix and approximating it by the first-order Chebyshev polynomial as a hypergraph convolution operation with single scale groupwise interaction dependence as output $\mathbf{Y}\in\mathbb{R}^{N\times d}$.
@@
Finally, we aggregate all the multiscale hypergraphs together to present crowded miscellaneous groupwise interactions as follows:
@@
where $f_{MLP}$ is a MLP neural network, and $\mathbf{\hat{Y}_{\rm{H}}}\in\mathbb{R}^{N\times H\times d}$ is the groupwise interaction feature.
@@ III-E Multi-Modal Fusion Network and CVAE Decoder
Subsequently, the groupwise HHI features $\mathbf{\hat{Y}}_{\rm{HS}},\mathbf{\hat{Y}}_{\rm{HT}}$ and pairwise HHI dependencies $\mathbf{\hat{Y}}_{\rm{ST}}$ are aligned by a multi-head self-attention network as final crowd dynamics representation $\mathbf{\hat{Y}}_{\rm{M}}$.
@@
Eventually, to estimate the stochasticity of human movements, a conditional variational auto encoder (CVAE)-based decoder [22] is employed to approximate maximum likelihood in potential distributions of motion uncertainty. The environmental dynamics feature $\mathbf{\hat{Y}}_{\rm M}$ and observed data $\mathbf{X}$ are encoded to present Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$}. And the latent variable $z$ is sampled by above Gaussian distribution $z\sim{\mathcal{N}(0,\sigma_{\rm T_{i}}^{2}\mathbf{I})}$ in the testing process. Lastly, the maximum likelihood forecasting result is calculated by the decoder based on the concatenation of latent variable and observation embedding in the CVAE block as follows:
@@
where $\mu,\sigma$ are the mean and variance of the approximate distribution, and $Encoder(\cdot)$ is Hyper-STTN backbone framework, and $Decoder(\cdot)$ is a spectral temporal graph network from [31].
@@ IV-A1 Dataset
We conduct experiments on several widely used public pedestrian trajectory datasets collected from real-world scenarios. The training and validation data comprise the ETH-UCY [32] and NBA [5] datasets, which record pedestrians’ temporal location information in world coordinates. Consistent with prior works, our experiments adopt a leave-one-out cross-validation pre-process method, reserving only the last subsection of each scene for testing.
@@ IV-A2 Evaluation Metrics
In our experiments, we evaluate prediction accuracy using the average displacement error (ADE) and final displacement error (FDE) metrics [4]. The ADE20/FDE20 quantify performance with the Euclidean distance between predicted and ground-truth trajectories, either averaged over all timesteps or at the final timestep, which evaluate performance on the top 20 sampled predictions.
@@ IV-B1 Baselines
We compared our model with several existing SOTA algorithms. The algorithms included in the comparison are Social-Attention [16], Social-GAN [6], Social-STGCNN [18], Trajectron++ [9], STAR [7], PECNet [33], Social-Implicit [19], GroupNet [5], EqMotion [34], LED [13], and IMLE [14], as shown in Table. I and Table. II.
@@ IV-B2 Ablation Models
We design two ablation models to evaluate the contributions of distinct components for overall performance. The ablation model STTN, infers pedestrian interactions using only the hybrid spatial-temporal transformer, excluding the hypergraph network block from Hyper-STTN. The second mode, HGNN, estimates crowd motion dependencies solely through hypergraph convolution operations without transformers. In both ablation models, the CVAE structure is retained as the decoder. All training configurations, including datasets and initialization parameters, are kept identical to those used for Hyper-STTN.
@@
Stochastic
ADE20 / FDE20
Social-
Attention [16]
ICRA18
Social-
GAN [6]
CVPR18
Trajectron
++ [9]
ECCV20
STAR
[7]
ECCV20
Social-
Implicit [19]
ECCV22
GroupNet
[5]
CVPR22
EqMotion
[34]
CVPR23
LED
[13]
CVPR23
IMLE
[14]
CVPR25
STTN
HGNN
Hyper-STTN
ETH 1.39/2.39 0.87/1.62 0.61/1.02 0.36/0.65 0.61/1.08 0.46/0.73 0.40/0.61 0.39/0.58 0.40/0.58 0.52/0.75 0.42/0.73 0.35/0.57
HOTEL 2.51/2.91 0.67/1.37 0.19/0.28 0.17/0.36 0.33/0.63 0.15/0.25 0.12/0.18 0.11/0.17 0.12/0.18 0.21/0.25 0.21/0.32 0.14/0.18
UNIV 1.25/2.54 0.76/1.52 0.30/0.54 0.31/0.62 0.52/1.11 0.26/0.49 0.23/0.43 0.26/0.44 0.23/0.39 0.25/0.43 0.30/0.53 0.23/0.41
ZARA1 1.01/2.17 0.35/0.68 0.24/0.42 0.26/0.55 0.32/0.66 0.21/0.39 0.18/0.32 0.18/0.26 0.16/0.26 0.31/0.39 0.28/0.53 0.18/0.28
ZARA2 0.88/1.75 0.42/0.84 0.18/0.32 0.22/0.46 0.43/0.85 0.17/0.33 0.13/0.23 0.13/0.22 0.13/0.22 0.25/0.33 0.23/0.28 0.15/0.22
AVG 1.41/2.35 0.61/1.21 0.30/0.51 0.26/0.53 0.33/0.67 0.25/0.44 0.21/0.35 0.21/0.33 0.21/0.33 0.31/0.43 0.29/0.48 0.21/0.33
TABLE II: The crowded trajectory forecasting best-of-20 stochastic sampled results of minADE20 / minFDE20 on NBA dataset.
Stochastic
ADE20 / FDE20
Social-GAN [6]
CVPR18
Social-STGCNN [18]
CVPR20
Trajectron++ [9]
ECCV20
PECNet [33]
ECCV20
GroupNet [5]
CVPR22
STTN
HGNN
Hyper-STTN
NBA-1s 0.41/0.62 0.34/0.48 0.30/0.38 0.35/0.58 0.34/0.48 0.39/0.58 0.37/0.59 0.30/0.37
NBA-2s 0.81/1.32 0.71/0.94 0.59/0.82 0.68/1.23 0.62/0.95 0.72/1.05 0.64/0.95 0.58/0.81
NBA-3s 1.19/1.94 1.09/1.77 0.85/1.24 1.01/1.76 0.87/1.31 1.07/1.51 0.98/1.44 0.84/1.21
NBA-4s 1.59/2.41 1.53/2.26 1.15/1.57 1.31/1.79 1.13/1.69 1.24/1.99 1.18/1.89 1.01/1.52
@@ IV-B3 Training Details
We use two NVIDIA RTX-4090 GPUs to train the model via the Adam optimizer with a decay learning rate of $1\times 10^{-4}$. To improve robustness, Gaussian noise is injected during training. Additional parameters and implementation details are available on the project website.
@@
where $\kappa$ is the weight of each loss term, and the Kullback-Leibler (KL) divergence term $f_{\mathrm{KL}}(\cdot||\cdot)$} is used to update the encoder and decoder in the CVAE block, and $^{\mathbb{{\hat{X}}}}$ denotes the groundturth of dataset. The $\angle(\cdot)$ function computes the angle between two vectors representing points.
@@
IV-C Quantitative Results
We conducted the trajectory prediction task with respect to stochastic conditions, performing Hyper-STTN’s SOTA quantitative results in ETH-UCY and NBA datasets, as shown in Table. I and Table. II. Notably, due to the datasets’ size and the rapid advancement of expressive models, the performance on the ETH-UCY datasets has plateaued in recent years that is approaching the inherent errors of machine learning approaches on such datasets. However, Hyper-STTN still exhibits comparable performance against existing advanced models both on ADE20 and FDE20 metrics. For instance, in the ETH-UCY dataset, Hyper-STTN not only enhances the $12.5\%$ ADE20 and $1.7\%$ FDE20 from EqMotion’s [34] 0.40/0.58 to 0.35/0.57 on ETH dataset, but also improves the average $14.7\%/6.5\%/3.4\%/10.6\%$ ADE20 and $22.9\%/14.7\%/7.6\%/10.1\%$ FDE20 performance from Groupnet’s [5] on NBA dataset. Highlighting the effectiveness of Hyper-STTN in crowd latent interaction inference and trajectory forecasting.
Deep learning–based trajectory prediction frameworks have catalyzed a shift toward modeling latent social interactions across spatial and temporal dimensions for pedestrian motion forecasting. Methods built upon recurrent architectures [4], attention mechanisms [16], and, more recently, transformer models [7, 8] have demonstrated progressively stronger predictive performance by enabling richer high-level feature representations. Following this trend, Hyper-STTN adopts transformer backbones to explicitly model fundamental social interactions and temporal dependencies in dense crowds. Its ablation variant, STTN, which is purely transformer-driven (akin to STAR[7, 8]), isolates the contribution of these components. Experimental results indicate that the observed performance gains arise from the introduction of masked-attention and cross-modal attention modules. These modules mitigate agent-flicker effects relative to vanilla self-attention under fixed, predefined transformer data scales, while simultaneously improving the adaptability of spatial–temporal fusion.
Despite the remarkable effectiveness of transformers in abstracting pairwise interactions, their limited consideration of group-level dynamics constrains further performance improvements—particularly in dense scenes where individual motions are strongly influenced by the movements of surrounding groups. For example, GroupNet [5] achieves superior performance (0.26/0.49) compared to STAR [7] (0.31/0.62) on the UNIV dataset, which contains more densely crowded scenarios. To address this gap, we propose to integrate hypergraph construction for inferring groupwise interactions with transformer architectures for pairwise feature modeling, thereby capturing richer strategies for crowded movement prediction. In addition, the ablation model HGNN exhibits performance comparable to GroupNet across several datasets, underscoring the effectiveness of the hypergraph component.